Operation Summary for October 14, 2009 - October 16, 2009 |
Wednesday Oct. 14 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (104.9E10), with the Antiproton Source stacking (44.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. During the day shift, Operations established HEP store 7265 at 8:30 AM with an average initial luminosity (AIL) of 243.5E30. Operators prepared the Booster, MI, MiniBooNE and NuMI enclosures for accesses. They were all out of access by 2:45 PM. A failed MI extraction magnet power supply held off Pbar. During the evening shift, MI personnel had the extraction magnet power supply fixed by 4:30 PM. Thursday Oct. 15 The midnight shift began with Operations monitoring store 7265 (62.24E30) and the stash (179.4E10), with Pbar stacking (25.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7265 at 12:30 AM. CDF accessed their collision hall until 2 AM. Operations put the TeV into shot setup and established store 7269 at 6 AM with an AIL of 229E30. During the day shift, Switchyard and MI personnel conducted MI-52 septa studies. During the evening shift, a MI-40 lower power supply failed at 9:54 PM. Operations called in power supply personnel who found a bad microswitch. Friday Oct. 16 The midnight shift began with Operations monitoring store 7269 (52.08E30), the stash (349.0E10), and the Pbar stack (18.1E10), with MI-40 power supply work continuing, and with NuMI and MiniBooNE taking beam. During the midnight shift, power supply technicians had the MI-40 power supply fixed by 12:15 AM. Operations terminated store 7269 at 1:15 AM. Operations established store 7270 with an AIL of 268E30 at 3 AM. At 7:49 AM, the TeV sector F1 cryo system went unstable and quenched. The Plans for Friday and the Weekend The plans for today are to stack, stash, and store. The Run Coordinator will allow CDF and D0 to access their collision halls and allow TeV personnel to conduct studies. Cryo system personnel will investigate the cause of the F1 quench. |
|
More Information Comments and Suggestions |
| last modified 10/16/2009 email Fermilab |
| Security, Privacy, Legal | |